OpenStudy (anonymous):

help! very desperate rn! If h(x)=(fog)(x) and h(x)=5(x+1)^3 find f(x) and g(x) thank you!

random231 (random231):

f(x)=x+1 and g(x)=5*x^3 ?

OpenStudy (anonymous):

thank you so much! @random231 would you be willing to help me with some more? i was dumb this semester and managed to fall very very far behind and today is my last day to turn in work

random231 (random231):

sure!

OpenStudy (anonymous):

for f(x)=2x+1 and g(x)=x^2-7 find (fog)(x)

random231 (random231):

fog(x)=2(g(x))+1

OpenStudy (anonymous):

2x^2-13?

random231 (random231):

bullseye!

OpenStudy (anonymous):

when it's (f/g)(x) what do you do?

random231 (random231):

its simply f(x)/g(x)

OpenStudy (anonymous):

oh okay! i think i'm getting the hang of it thank you :)

random231 (random231):

no problem

OpenStudy (anonymous):

in another one it's (fog)(2) do i substitute x for 2 and do the same as before?

random231 (random231):

find fog(x) then substitute x with 2 and find the value

OpenStudy (anonymous):

oh okay! i got -5